write an equation for the problems and solve : Mary and her father took an all day automobile ride, going 3 times as far in the afternoon as in the morning. In all they traveled 420 miles. How far did they travel in the morning?
Answer by Theo(13342)

let x represent the distance they traveled in the morning.
then the distance they traveled in the afternoon is 3x.
the total distance they traveled is 420 miles.
this means that 3x + x = 420 which becomes 4x = 420
divide both sides of this equation by 4 to get:
x = 420/4 = 105.
they traveled 105 miles in the morning.
they traveled 3 * 105 miles in the afternoon = 315 miles.
total distance they traveled is 105 + 315 = 420 miles.
answer to the problem is that they traveled 105 miles in the morning.